# Phi-3 Technical Report: A Highly Capable Language Model Locally on Your Phone

### v1

@@ 2 Technical Specifications
Our models went through post-training with both supervised instruction fine-tuning, and preference tuning with DPO. We have worked on generating and curating various instruction and preference data. This has improved the model chat capabilities, robustness, as well as its safety.
@@ 3 Academic benchmarks
Phi-3-mini 3.8b Phi-3-small 7b Phi-3-medium 14b (preview) Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Mixtral 8x7b GPT-3.5 version 1106 MMLU (5-Shot) [HBK+21] 68.8 75.3 78.2 56.3 61.7 63.6 66.0 68.4 71.4 HellaSwag (5-Shot) [ZHB+19] 76.7 78.7 83.0 53.6 58.5 49.8 69.5 70.4 78.8 ANLI (7-Shot) [NWD+20] 52.8 55.0 58.7 42.5 47.1 48.7 54.8 55.2 58.1 GSM-8K (0-Shot; CoT) [CKB+21] 82.5 88.9 90.3 61.1 46.4 59.8 77.4 64.7 78.1 MedQA (2-Shot) [JPO+20] 53.8 58.2 69.4 40.9 49.6 50.0 58.9 62.2 63.4 AGIEval (0-Shot) [ZCG+23] 37.5 45.0 48.4 29.8 35.1 42.1 42.0 45.2 48.4 TriviaQA (5-Shot) [JCWZ17] 64.0 59.1 75.6 45.2 72.3 75.2 73.6 82.2 85.8 Arc-C (10-Shot) [CCE+18] 84.9 90.7 91.0 75.9 78.6 78.3 80.5 87.3 87.4 Arc-E (10-Shot) [CCE+18] 94.6 97.1 97.8 88.5 90.6 91.4 92.3 95.6 96.3 PIQA (5-Shot) [BZGC19] 84.2 87.8 87.7 60.2 77.7 78.1 77.1 86.0 86.6 SociQA (5-Shot) [BZGC19] 76.6 79.0 80.2 68.3 74.6 65.5 73.2 75.9 68.3 BigBench-Hard (0-Shot) [SRR+22, SSS+22] 71.7 75.0 81.3 59.4 57.3 59.6 68.9 69.7 68.32 WinoGrande (5-Shot) [SLBBC19] 70.8 82.5 81.4 54.7 54.2 55.6 58.0 62.0 68.8 OpenBookQA (10-Shot) [MCKS18] 83.2 88.4 87.2 73.6 79.8 78.6 81.6 85.8 86.0 BoolQ (0-Shot) [CLC+19] 77.2 82.9 86.6 72.2 66.0 78.3 77.6 79.1 CommonSenseQA (10-Shot) [THLB19] 80.2 80.3 82.6 69.3 72.6 76.2 73.6 78.1 79.6 TruthfulQA (10-Shot) [LHE22] 65.0 68.7 75.7 52.1 53.0 62.0 60.1 85.8 HumanEval (0-Shot) [CTJ+21] 58.5 59.1 55.5 59.0 28.0 34.1 38.4 37.8 62.2 MBPP (3-Shot) [AON+21] 70.0 71.4 74.5 60.6 50.8 51.5 65.3 60.2 77.8 Average 71.2 74.9 78.2 61.0 62.0 68.0 69.9 75.3 GPQA (2-Shot; CoT) [RHS+23] 32.8 34.3 29.0 MT Bench (2 round ave.) [ZCS+23] 8.38 8.70 8.91 8.35
@@ 4 Safety
Table 1 shows the results of in-house RAI benchmarks for phi-3-mini-4k and phi-3-mini-128k compared to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Gemma 7b [TMH+24]. This benchmark utilized GPT-4 to simulate multi-turn conversations in five different categories and to evaluate the model responses. Ungroundedness between 0 (fully grounded) and 4 (not grounded) measures if the information in a response is based on a given prompt. In other categories, responses were evaluated in terms of the severity of harmfulness from 0 (no harm) to 7 (extreme harm) and the defect rates (DR-$x$) were computed as the percentage of samples with the severity score being greater than or equal to $x$.
@@ Appendix B Authors
Marah Abdin Russell J. Hewett Corby Rosset
Sam Ade Jacobs Jamie Huynh Olatunji Ruwase
Ammar Ahmad Awan Mojan Javaheripi Olli Saarikivi
Jyoti Aneja Xin Jin Amin Saied
Ahmed Awadallah Piero Kauffmann Adil Salim
Hany Awadalla Nikos Karampatziakis Michael Santacroce
Nguyen Bach Dongwoo Kim Shital Shah
Amit Bahree Mahoud Khademi Ning Shang
Arash Bakhtiari Lev Kurilenko Hiteshi Sharma
Harkirat Behl James R. Lee Xia Song
@@
Abhishek Goswami Thomas Portet Li Lyna Zhang
Suriya Gunasekar Reid Pryzant Yi Zhang
Emman Haider Sambudha Roy Yunan Zhang
Junheng Hao Marko Radmilac Xiren Zhou

### v2

@@ 2 Technical Specifications
Post-training of phi-3-mini went through two stages, including supervised finetuning (SFT) and direct preference optimization (DPO). SFT leverages highly curated high-quality data across diverse domains, e.g., math, coding, reasoning, conversation, model identity, and safety. The SFT data mix starts with using English-only examples. DPO data covers chat format data, reasoning, and responsible AI (RAI) efforts. We use DPO to steer the model away from unwanted behavior, by using those outputs as “rejected” responses. Besides improvement in math, coding, reasoning, robustness, and safety, post-training transforms a language model to an AI assistant that users can efficiently and safely interact with.
As part of the post-training process, we developed a long context version of phi-3-mini with context length limit enlarged to 128K instead of 4K. Across the board, the 128K model quality is on par with the 4K length version, while being able to handle long context tasks. Long context extension has been done in two stages, including long context mid-training and long-short mixed post-training with both SFT and DPO.
@@ 3 Academic benchmarks
Phi-3-mini 3.8b Phi-3-small 7b (preview) Phi-3-medium 14b (preview) Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Mixtral 8x7b GPT-3.5 version 1106 MMLU (5-Shot) [HBK+21] 68.8 75.3 78.2 56.3 61.7 63.6 66.0 68.4 71.4 HellaSwag (5-Shot) [ZHB+19] 76.7 78.7 83.0 53.6 58.5 49.8 69.5 70.4 78.8 ANLI (7-Shot) [NWD+20] 52.8 55.0 58.7 42.5 47.1 48.7 54.8 55.2 58.1 GSM-8K (0-Shot; CoT) [CKB+21] 82.5 88.9 90.3 61.1 46.4 59.8 77.4 64.7 78.1 MedQA (2-Shot) [JPO+20] 53.8 58.2 69.4 40.9 49.6 50.0 58.9 62.2 63.4 AGIEval (0-Shot) [ZCG+23] 37.5 45.0 48.4 29.8 35.1 42.1 42.0 45.2 48.4 TriviaQA (5-Shot) [JCWZ17] 64.0 59.1 75.6 45.2 72.3 75.2 73.6 82.2 85.8 Arc-C (10-Shot) [CCE+18] 84.9 90.7 91.0 75.9 78.6 78.3 80.5 87.3 87.4 Arc-E (10-Shot) [CCE+18] 94.6 97.1 97.8 88.5 90.6 91.4 92.3 95.6 96.3 PIQA (5-Shot) [BZGC19] 84.2 87.8 87.7 60.2 77.7 78.1 77.1 86.0 86.6 SociQA (5-Shot) [BZGC19] 76.6 79.0 80.2 68.3 74.6 65.5 73.2 75.9 68.3 BigBench-Hard (0-Shot) [SRR+22, SSS+22] 71.7 75.0 81.3 59.4 57.3 59.6 68.9 69.7 68.32 WinoGrande (5-Shot) [SLBBC19] 70.8 82.5 81.4 54.7 54.2 55.6 58.0 62.0 68.8 OpenBookQA (10-Shot) [MCKS18] 83.2 88.4 87.2 73.6 79.8 78.6 81.6 85.8 86.0 BoolQ (0-Shot) [CLC+19] 77.2 82.9 86.6 72.2 66.0 78.3 77.6 79.1 CommonSenseQA (10-Shot) [THLB19] 80.2 80.3 82.6 69.3 72.6 76.2 73.6 78.1 79.6 TruthfulQA (10-Shot) [LHE22] 65.0 68.7 75.7 52.1 53.0 62.0 60.1 85.8 HumanEval (0-Shot) [CTJ+21] 59.1 59.1 55.5 47.0 28.0 34.1 60.4 37.8 62.2 MBPP (3-Shot) [AON+21] 70.0 71.4 74.5 60.6 50.8 51.5 65.3 60.2 77.8 Average 71.2 74.9 78.2 61.0 62.0 68.0 69.9 75.3 GPQA (2-Shot; CoT) [RHS+23] 32.8 34.3 29.0 MT Bench (2 round ave.) [ZCS+23] 8.38 8.70 8.91 8.35
@@ 4 Safety
Table 1 shows the results of in-house RAI benchmarks for phi-3-mini-4k and phi-3-mini-128k compared to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Gemma 7b [TMH+24], and Llama-3-instruct-8b [AI23]. This benchmark utilized GPT-4 to simulate multi-turn conversations in five different categories and to evaluate the model responses. Ungroundedness between 0 (fully grounded) and 4 (not grounded) measures if the information in a response is based on a given prompt. In other categories, responses were evaluated in terms of the severity of harmfulness from 0 (no harm) to 7 (extreme harm) and the defect rates (DR-$x$) were computed as the percentage of samples with the severity score being greater than or equal to $x$.
@@ Appendix B Authors
Marah Abdin Russell J. Hewett Olatunji Ruwase
Sam Ade Jacobs Jamie Huynh Olli Saarikivi
Ammar Ahmad Awan Mojan Javaheripi Amin Saied
Jyoti Aneja Xin Jin Adil Salim
Ahmed Awadallah Piero Kauffmann Michael Santacroce
Hany Awadalla Nikos Karampatziakis Shital Shah
Nguyen Bach Dongwoo Kim Ning Shang
Amit Bahree Mahmoud Khademi Hiteshi Sharma
Arash Bakhtiari Lev Kurilenko Xia Song
Harkirat Behl James R. Lee Masahiro Tanaka
@@
Amit Garg Thomas Portet Li Lyna Zhang
Abhishek Goswami Reid Pryzant Yi Zhang
Suriya Gunasekar Sambuddha Roy Yue Zhang
Emman Haider Marko Radmilac Yunan Zhang
Junheng Hao Corby Rosset Xiren Zhou